OpenStudy (anonymous):

find the inverse of the one-to-one function 7x+5

OpenStudy (mathstudent55):

The given function is: y = 7x + 5 To find its inverse, follow these steps: 1. Switch x and y 2. Solve for y The new y function is the inverse of the original y function.

OpenStudy (anonymous):

so i got x=y+5/7 is that right?

OpenStudy (mathstudent55):

y = 7x + 5 Switch x and y: x = 7y + 5 Switch sides: 7y + 5 = x Subtract 5 from both sides: 7y = x - 5 Divide both sides by 7: y = (x - 5)/7 y = (1/7)x - 5/7 \(y = \dfrac{1}{7}x - \dfrac{5}{7} \)

OpenStudy (mathstudent55):

That is how you do it. Notice that the 5 is subtracted from y, not added.

OpenStudy (anonymous):

word